OpenStudy (anonymous):

Solve log(3x + 2) = 3.

OpenStudy (eyust707):

" log(3x + 2) = 3." is the same as saying" \[10^{3} = 3x + 2\]

OpenStudy (eyust707):

Do you know how to solve from there?

OpenStudy (anonymous):

Promote both sides to powers of 10: \(10^{\log(3x+2)}=10^3\). Since we know \(10^{\log{x}}=x\), we can simplify the left-hand side \(3x+2=10^3=1000\), which should seem rather simple to solve for \(x\).

OpenStudy (anonymous):

thanks!